OpenStudy (anonymous):

find general soln to: dy/dx=e^(2y-1)

OpenStudy (anonymous):

Separate the variables. Then integrate.

OpenStudy (anonymous):

yeah i got ln something.. should there be mod inside the ln ?

OpenStudy (anonymous):

We dont get ln.

OpenStudy (anonymous):

no after solving for y u do

OpenStudy (anonymous):

dx/dy=e^-2y+1 .: x=-1/2 e^-2y+1 +c -2(x-c)=e^-2y+1 -2y+1=ln(-2(x-c)) or should it be ln|-2(x-c)| ? mod or no mod inside ln

OpenStudy (anonymous):

@Abhishek619 will help.

OpenStudy (anonymous):

 now try to integrate on both sides.